# EUvsDisinfo: a Dataset for Multilingual Detection of Pro-Kremlin Disinformation in News Articles

### v1

@@ Abstract.
This work introduces EUvsDisinfo, a multilingual dataset of trustworthy and disinformation articles related to pro-Kremlin themes. It is sourced directly from the debunk articles written by experts leading the EUvsDisinfo project. Our dataset is the largest to-date resource in terms of the overall number of articles and distinct languages. It also provides the largest topical and temporal coverage. Using this dataset, we investigate the dissemination of pro-Kremlin disinformation across different languages, uncovering language-specific patterns targeting specific disinformation topics. We further analyse the evolution of topic distribution over an eight-year period, noting a significant surge in disinformation content before the full-scale invasion of Ukraine in 2022. Lastly, we demonstrate the dataset’s applicability in training models to effectively distinguish between disinformation and trustworthy content in multilingual settings.
@@ 1. Introduction
Combating pro-Kremlin disinformation articles presents significant challenges due to its dissemination across various languages, during an extensive period of time, and covering a wide range of different topics. Current disinformation datasets are not suitable for training the models aimed at this task since they lack the combination of such characteristics. In particular, the majority of such datasets focus on analysing short claims (Thorne et al., 2018; Wang, 2017) or social media posts (Potthast et al., 2018; Zubiaga et al., 2016). Out of the existing article-level disinformation corpora, the majority consist of monolingual data (Hossain et al., 2020; Shu et al., 2020; Karadzhov et al., 2017; Hardalov et al., 2016; Wang et al., 2020; Golbeck et al., 2018; Pérez-Rosas et al., 2018; Shu et al., 2020; Shahi et al., 2021). A significantly smaller amount of multilingual datasets is available (De et al., 2021; Dementieva et al., 2023; Li et al., 2020; Kishore Shahi and Nandini, 2020), and they contain the following limitations: (i) representing parallel translated data (De et al., 2021; Dementieva et al., 2023), (ii) covering short periods of time ( 1 year) (De et al., 2021; Li et al., 2020; Kishore Shahi and Nandini, 2020), (iii) containing few articles (around ) (De et al., 2021), (iv) focusing on narrow topics (De et al., 2021; Li et al., 2020; Kishore Shahi and Nandini, 2020). Furthermore, to the best of our knowledge, the work by Solopova et al. (2023) is the only dataset addressing the theme of pro-Kremlin information warfare, although their dataset is targeted towards the detection of a pro-Western vs pro-Kremlin stance rather than disinformation.
@@
In this work, we leverage the journalistic investigation carried out by specialists to generate a large and diverse multilingual dataset containing trustworthy and disinformation articles related to pro-Kremlin disinformation narratives. To produce such a dataset, we use the debunk articles written by EUvsDisinfo111https://euvsdisinfo.eu, who since 2015 has been providing an EU-wide response to information warfare by debunking disinformation narratives across EU-member states. Our dataset is the largest ( articles) to-date, most topically diverse ( topics), spans over the longest period of time ( years), and is most diverse in terms of languages ( languages) compared to other article-level multilingual disinformation detection datasets. Table 1 provides an overview of related datasets in comparison to ours. Our contributions can be summarised as follows: We (i) provide 222Dataset: https://doi.org/10.5281/zenodo.10514307 the largest and most diverse article-level multilingual disinformation dataset in terms of language, topics, and time periods. (ii) analyse the prominence of pro-Kremlin disinformation topics across different languages and time periods. (iii) evaluate several baseline models using our novel dataset for the task of multilingual binary disinformation detection. Our dataset and code are made publicly available, along with documentation, as supplementary material to facilitate reproducibility333Supplementary material: https://github.com/JAugusto97/euvsdisinfo.
@@ 2. Methodology
Our dataset leverages the structured information of the debunking articles published by EUvsDisinfo (see Figure 1 for an example). It consists of two classes of articles: disinformation and trustworthy. The disinformation articles are acquired directly from the links mentioned on the left hand-side of the EUvsDisinfo debunk page (see Figure 1). The trustworthy articles are derived from the links to reliable sources within the response section of the article, in which EUvsDisinfo directly debunks the false narrative. Such URLs can lead to news articles, official documents, statements, books, encyclopedia articles, social media posts, and fact-checking articles, among others.
To verify that EUvsDisinfo does not reference disinformation articles in the response section, we manually inspected a randomly sampled set of 30 debunk articles, and annotated all the URLs mentioned within their response section. In total, 350 URLs were labelled as either trustworthy or potential disinformation. To do so, we considered the context in which the URL is referenced in the response section. For example, if the URL is referenced after a sentence such as ”see similar cases”, it is marked as potential disinformation (for reference, see the last paragraph in the response section of Figure 1). As a result of this sanity check, we identified that all of the URLs marked as potential disinformation actually correspond to other debunk articles by EUvsDisinfo or other fact-checking agencies such as Bellingcat and StopFake, thus none correspond to disinformation.
@@ 2.1. Data Collection
Due to the wide variety of websites mentioned in EUvsDisinfo debunks, extracting the textual content out of their varying HTML structure is non-trivial. To facilitate this task, we employ the Diffbot API444https://www.diffbot.com, which is a proprietary tool that uses machine learning to extract the content of a given web page. Although Diffbot is a closed-source service, it is free of charge for academic purposes. For articles that are no longer available on the web, we use the digital library Wayback Machine555https://web.archive.org to retrieve the latest achieved version of the web page. Also, the language of each trustworthy article is inferred using Polyglot666https://github.com/aboSamoor/polyglot, as EUvsDisinfo only specifies the languages of the disinformation articles.
Upon collecting the content of the URLs, we apply three filtering strategies to ensure consistency in the dataset. First, we remove URLs whose domain is not of a news outlet type. This is done to ensure consistency in the theme of disinformation detection of news articles published by news outlets. To do so, we first manually inspect the most frequent domains of the URLs, and label them as one of the following: “News Outlet”, “Organisation”, “Social Media”, “Fact-checker”, and “Other”. Next, we remove instances referring to error messages, log-in prompts, and paywalls through lexicon-bases rules, and through removing articles with less than characters, which is significantly shorter than the average article ( characters). Lastly, we remove URLs cited within sentences containing n-grams referring to other fact-checking articles (e.g. ”See earlier disinformation cases”). The full list of n-grams can be found in the supplementary materials.
@@ 2.2. Dataset Overview
The full dataset is comprised of news articles with () and () marked as disinformation and trustworthy, respectively. It contains articles by different publishers on unique topics, spanning across years, between 06/01/2015 and 01/08/2023. On average, the articles are characters in length and cover different topics. The dataset contains different languages, of which are not present on existing multilingual article-level disinformation datasets presented in Table 1. Please refer to the supplementary material for a detailed breakdown of all languages. Out of the total of publishers in the dataset, and are associated with disinformation and trustworthy articles, respectively. The top 5 publishers of disinformation articles are Russian outlets: Sputnik ( ), RT (), RIA Novosti (), Tsargrad (), and Ukraina.ru (). For trustworthy articles, the top 5 publishers include Reuters (), BBC (), The Guardian (), Deutsche Welle (), and Radio Free Europe ().
@@ 2.3. Analysis of Topics
Figure 2 shows the distribution of the eight most frequent topics across the whole dataset for the five most common languages. Notably, the topic of “War in Ukraine” is the most prevalent across all five languages. However, certain topics appear more frequently in specific languages. The topic “Invasion of Ukraine” is over twice more common in Spanish compared to other languages. Moreover, “Conspiracy theory” and “Crimea” are frequently discussed in four out of the top five languages, except Spanish, where they account for only and of articles respectively, compared to an average of and in other languages. A similar discrepancy is noted for ”Invasion of Ukraine” in German, appearing in just of articles, while averaging in Russian, Spanish, Arabic, and English. These disparities suggest a language-specific targeted dissemination of disinformation contingent on the topics.
Next, to examine the emergence of related themes over time, we group similar high-frequency topics into four overarching themes: “COVID-19”, “West”, “Russia”, and “Ukraine”. These themes are manually defined by merging similar topics from the top 50 most frequent in the dataset. Figure 3 shows the temporal distribution of disinformation topics under each theme. From the first quarter of 2021 onwards, there is a noticeable increase in disinformation articles across all themes, coinciding with Russia’s escalating military presence along the Ukrainian border. Similar surge in the number of “COVID-19”-related topics can be observed in the months following the start of the pandemic. Following the full-scale invasion of Ukraine in the first quarter of 2022, disinformation articles on all topics decline significantly, except for ”War in Ukraine” and ”Invasion of Ukraine,” which become the dominant disinformation topics.
@@ 3. Experimental Setup
We employ the EUvsDisinfo dataset to perform binary disinformation detection (i.e. classify news articles as either disinformation or trustworthy). While the full dataset contains articles in unique languages, several languages are severely underrepresented and/or highly skewed towards one of the classes. Therefore in order to draw reliable conclusions, we discard underrepresented languages (fewer than articles) and those in which more than 95% of the articles belong to only one of the two classes. The resulting dataset used in the experiments comprises articles distributed across languages (% of disinformation articles is also shown): English (, ), Russian (, ), German (, ), French (, ), Spanish (, ), Georgian (, ), Czech (, ), Polish (, ), Italian (, ), Lithuanian (, ), Romanian (, ), Slovak (, ), Serbian (, ), Finnish (, ).
We experiment with four models that have been used in prior work on article-level disinformation classification. As baselines, we train a Multinomial Naive Bayes (MNB), and a Support Vector Machine (SVM), both using bag-of-words to encode the textual features. MNB and SVM are classic machine learning baselines commonly adopted in other disinformation detection works (Shu et al., 2020; Golbeck et al., 2018; Li et al., 2020; Solopova et al., 2023). We also train transformer-based models, mBERT and XLM-RoBERTa, as three of the relevant multilingual datasets discussed in Section 1 use mBERT as a baseline (De et al., 2021; Solopova et al., 2023; Kishore Shahi and Nandini, 2020), and XLM-R is used by the fourth dataset Li et al. (2020).
We conduct a hyperparameter search for each method in the following manner: we first split the dataset into train and development sets comprising and of the data, respectively. The highest scoring hyperparameter configuration with respect to the development set is used throughout the experiments. Next, the development set is discarded, and the remaining of the dataset is split into train and test sets using a leave-one-out 10-fold cross-validation strategy. The folds are stratified by both language and class. Model performance for each language is measured using the score to account for the skewed distribution of classes. To obtain a unified metric that encapsulates the overall system performance across all languages, we further average the scores per language to obtain the score. Further details such as hyperparameters and hardware configurations can be found in the supplementary material.
@@ 5. Conclusions and Future Work
This paper introduced EUvsDisinfo, a large, linguistically, temporally, and topically diverse dataset of pro-Kremlin disinformation articles. Using the dataset we found evidences of language-specific targeting of specific topics, and revealed a surge in disinformation content related to the war in Ukraine after its full-scale invasion in 2022. Lastly, we proposed classification baselines using our dataset for the task of binary disinformation detection in a multilingual setting. In future work, we plan to leverage the structure of our dataset to explore evidence-aware fact-checking approaches by linking disinformation and trustworthy articles referring to the same narrative.
@@
The dataset is seeded from the publicly available debunks published by EUvsDisinfo. The content originates from news articles, thus it does not contain personal data. We recognise that the dataset is susceptible to misuse for malicious purposes, and we strongly urge researchers to use it in accordance with best practice ethics protocols. Our dataset is compliant with FAIR principles (Wilkinson et al., 2016). It is made fully available with a unique digital object identifier, in a CSV format that can be processed by most widely used tools, and is released under an Apache 2.0 license. Since the textual content of news articles may be copyrighted, we do not include them in the dataset, but instead provide a software (https://doi.org/10.5281/zenodo.10492913) to allow researchers to collect the content themselves.

### v3

@@ Abstract.
This work introduces EUvsDisinfo, a multilingual dataset of disinformation articles originating from pro-Kremlin outlets, along with trustworthy articles from credible / less biased sources. It is sourced directly from the debunk articles written by experts leading the EUvsDisinfo project. Our dataset is the largest to-date resource in terms of the overall number of articles and distinct languages. It also provides the largest topical and temporal coverage. Using this dataset, we investigate the dissemination of pro-Kremlin disinformation across different languages, uncovering language-specific patterns targeting certain disinformation topics. We further analyse the evolution of topic distribution over an eight-year period, noting a significant surge in disinformation content before the full-scale invasion of Ukraine in 2022. Lastly, we demonstrate the dataset’s applicability in training models to effectively distinguish between disinformation and trustworthy content in multilingual settings.
@@ 1. Introduction
Combating pro-Kremlin disinformation articles presents significant challenges due to its dissemination across various languages, during an extensive period of time, and covering a wide range of different topics. Current disinformation datasets are not suitable for training machine learning models aimed to address this task since they lack the combination of such characteristics. Most datasets focus on analysing short claims (Thorne et al., 2018; Wang, 2017) or social media posts (Potthast et al., 2018; Zubiaga et al., 2016), whilst the majority of the existing article-level disinformation corpora consist of monolingual data (Hossain et al., 2020; Shu et al., 2020; Karadzhov et al., 2017; Hardalov et al., 2016; Wang et al., 2020; Golbeck et al., 2018; Pérez-Rosas et al., 2018; Shu et al., 2020; Shahi et al., 2021). A significantly smaller amount of multilingual datasets is available (De et al., 2021; Dementieva et al., 2023; Li et al., 2020; Kishore Shahi and Nandini, 2020), and they contain the following limitations: (i) representing parallel translated data (De et al., 2021; Dementieva et al., 2023), (ii) covering short periods of time ( 1 year) (De et al., 2021; Li et al., 2020; Kishore Shahi and Nandini, 2020), (iii) containing only few articles (around ) (De et al., 2021), and (iv) focusing on narrow topics (De et al., 2021; Li et al., 2020; Kishore Shahi and Nandini, 2020). To the best of our knowledge, the work by Solopova et al. (2023) is the only dataset addressing the theme of pro-Kremlin information warfare. Nevertheless, their dataset is targeted towards the detection of a pro-Western vs pro-Kremlin stance rather than disinformation.
@@
In this work, we leverage the journalistic investigation carried out by specialists to generate a large and diverse multilingual dataset containing disinformation articles from pro-Kremlin outlets and trustworthy counterparts from reliable / less biased sources. To produce such a dataset, we use the debunk articles written by EUvsDisinfo111https://euvsdisinfo.eu, who since 2015 has been providing an EU-wide response to information warfare by debunking disinformation narratives across EU-member states. Our dataset is the largest ( articles) to-date, most topically diverse ( topics manually assigned by EUvsDisinfo), spans over the longest period of time ( years), and is the most diverse in terms of languages ( languages) compared to other article-level multilingual disinformation detection datasets (see Table 1 for an overview of related datasets in comparison to ours). Our key contributions are: (i) The novel multilingual disinformation dataset222Dataset: https://doi.org/10.5281/zenodo.10514307, becoming the largest and most diverse article-level set in terms of language, topics, and time periods. (ii) The analysis of the prominence of pro-Kremlin disinformation topics across different languages and time periods. (iii) The evaluation of several baseline models using our novel dataset for the task of multilingual binary disinformation detection. Our dataset and code are made publicly available, along with documentation, as supplementary material to facilitate reproducibility333Supplementary material: https://github.com/JAugusto97/euvsdisinfo.
@@ 2. Methodology
Our dataset leverages the structured information of the debunking articles published by EUvsDisinfo (see Figure 1 for an example). It consists of two classes of articles: disinformation and trustworthy. Disinformation articles are acquired directly from the links mentioned on the left hand-side of the EUvsDisinfo debunk page. Trustworthy articles are derived from the links to reliable sources within the response section of the article, in which EUvsDisinfo directly debunks the false narrative. Such URLs can lead to news articles, official documents, statements, books, encyclopedia articles, social media posts, and fact-checking articles, among others.
To verify that EUvsDisinfo does not reference disinformation articles in the response section, we manually inspected a randomly sampled set of 30 debunk articles, and annotated all the URLs mentioned within their response section. In total, 350 URLs were labelled as either trustworthy or potential disinformation. To do so, we considered the context in which the URL is referenced in the response section. For example, if the URL is referenced after a sentence such as ”see similar cases”, it is marked as potential disinformation (for reference, see the last paragraph in the response section of Figure 1). We identified that all of the URLs marked as potential disinformation actually correspond to other debunk articles by EUvsDisinfo or other fact-checking agencies such as Bellingcat and StopFake, instead of directly links to disinformation.
@@ 2.1. Data Collection
Given the wide variety of websites mentioned in EUvsDisinfo debunks, with varying HTML structure, extracting textual content is not trivial. For this task, we employ the Diffbot API444https://www.diffbot.com, which is a proprietary tool that uses machine learning to extract the content of a given web page. Although Diffbot is a closed-source service, it is free of charge for academic purposes. We retrieve articles that are no longer available on the web, by using the digital library Wayback Machine555https://web.archive.org. Also, the language of each trustworthy article is inferred using Polyglot666https://github.com/aboSamoor/polyglot, as EUvsDisinfo only specifies the languages of the disinformation articles.
Upon collecting the contents of all the URLs mentioned on EUvsDisinfo debunks, resulting in articles, we apply three filtering strategies to ensure consistency in the dataset. First, we remove URLs whose domain is not of a news outlet type. This is done to ensure consistency in the theme of disinformation detection of news articles published by news outlets. To do so, we first manually inspect the most frequent domains of the URLs, and label them as one of the following: “News Outlet”, “Organisation”, “Social Media”, “Fact-checker”, and “Other”. Next, we remove instances referring to error messages, log-in prompts, and paywalls through lexicon-bases rules, and through removing articles with less than characters, which is significantly shorter than the average article ( characters). Lastly, we remove URLs cited within sentences containing n-grams referring to other fact-checking articles (e.g. ”See earlier disinformation cases”). The full list of n-grams can be found in the supplementary materials\@footnotemark.
@@ 2.2. Dataset Overview
The full dataset has news articles with () and () marked as disinformation and trustworthy, respectively. It contains articles by different publishers on unique topics, spanning across years, between 06/01/2015 and 01/08/2023. On average, the articles are characters in length and cover different topics. The dataset contains different languages, of which are not present on existing multilingual article-level disinformation datasets presented in Table 1. Please refer to the supplementary material for a detailed breakdown of all languages. Out of all publishers in the dataset, and are associated with disinformation and trustworthy articles, respectively. The top five publishers of disinformation articles are known pro-Kremlin outlets: Sputnik ( ), RT (), RIA Novosti (), Tsargrad (), and Ukraina.ru (). For trustworthy articles, the top five publishers include Reuters (), BBC (), The Guardian (), Deutsche Welle (), and Radio Free Europe ().
@@ 2.3. Analysis of Topics
Figure 2 shows the distribution of the eight most frequent topics across the whole dataset for the five most common languages. Notably, the topic of “War in Ukraine” is the most prevalent across all five languages. However, certain topics appear more frequently in specific languages. The topic “Invasion of Ukraine” is over twice more common in Spanish than in other languages. “Conspiracy theory” and “Crimea” are frequently discussed in four out of the top five languages, except Spanish, where they account for only and of articles respectively, compared to an average of and in other languages. A similar discrepancy is noted for ”Invasion of Ukraine” in German, appearing in just of articles, while averaging in Russian, Spanish, Arabic, and English. These disparities suggest a language-specific targeted dissemination of disinformation contingent on the topics.
To examine the emergence of related themes over time, we group similar high-frequency topics into four overarching themes: “COVID-19”, “West”, “Russia”, and “Ukraine”. These themes are manually defined by merging similar topics from the top 50 most frequent in the dataset. Figure 3 shows the temporal distribution of disinformation topics under each theme. From the first quarter of 2021 onwards, there is a noticeable increase in disinformation articles across all themes, coinciding with Russia’s escalating military presence along the Ukrainian border. Similar surge in the number of “COVID-19”-related topics can be observed in the months following the start of the pandemic. After the full-scale invasion of Ukraine in the first quarter of 2022, disinformation articles on all topics decline significantly, except for ”War in Ukraine” and ”Invasion of Ukraine”, which become the dominant disinformation topics.
@@ 3. Experimental Setup
We employ the EUvsDisinfo dataset to perform binary disinformation detection (i.e. classify news articles as either disinformation or trustworthy). While the full dataset contains articles in unique languages, several languages are severely underrepresented and/or highly skewed towards one of the classes. Therefore in order to draw reliable conclusions, we discard underrepresented languages (fewer than articles) and those in which more than 95% of the articles belong to only one of the two classes. The resulting dataset used in the experiments comprises articles distributed across languages (% of disinformation articles is also shown)\@footnotemark: English (, ), Russian (, ), German (, ), French (, ), Spanish (, ), Georgian (, ), Czech (, ), Polish (, ), Italian (, ), Lithuanian (, ), Romanian (, ), Slovak (, ), Serbian (, ), Finnish (, ).
We experiment with four models that have been used in prior work on article-level disinformation classification. We train commonly adopted baselines from previous work (Shu et al., 2020; Golbeck et al., 2018; Li et al., 2020; Solopova et al., 2023): Multinomial Naive Bayes (MNB), and Support Vector Machine (SVM), both using bag-of-words to encode the textual features. We also train transformer-based models, mBERT and XLM-RoBERTa, as three of the relevant multilingual datasets discussed in Section 1 use mBERT as a baseline (De et al., 2021; Solopova et al., 2023; Kishore Shahi and Nandini, 2020), and XLM-R is used by Li et al. (2020). We conduct a hyperparameter search for each model by: (i) splitting the dataset into train and development sets with and of the data, respectively. The highest scoring hyperparameter configuration with respect to the development set is used throughout the experiments. (ii) discarding the development set, and splitting the remaining of the dataset into train and test sets using a 10-fold cross-validation strategy. The folds are stratified by both language and class. Model performance for each language is measured using the score to account for the skewed distribution of classes. To obtain a unified metric that encapsulates the overall system performance across all languages, we further average the scores per language to obtain the score. Further details such as hyperparameters and hardware configurations can be found in the supplementary material\@footnotemark.
@@ 5. Conclusions and Future Work
This paper introduced EUvsDisinfo, a large, linguistically, temporally, and topically diverse dataset of disinformation and trustworthy articles originating from pro-Kremlin and reliable / less biased outlets, respectively. Using the dataset, we found evidences of language-specific targeting of specific topics, and revealed a surge in disinformation content related to the war in Ukraine right before its full-scale invasion in 2022. Lastly, we proposed classification baselines using our dataset for the task of binary disinformation detection in a multilingual setting. In future work, we plan to leverage the structure of our dataset to explore evidence-aware fact-checking approaches by linking disinformation and trustworthy articles referring to the same narrative.
@@
The dataset is seeded from the publicly available debunks published by EUvsDisinfo. The authors of the manuscript are not part of the EUvsDisinfo organisation. The dataset content originates from news articles, thus it does not contain personal data. The dataset specifically targets pro-Kremlin disinformation; while pro-Western disinformation is not within the scope of this work. We recognise that the dataset is susceptible to misuse for malicious purposes (e.g., used by originators of disinformation to improve their techniques), and we strongly urge researchers to use it in accordance with best practice ethics protocols. Our dataset is compliant with FAIR principles (Wilkinson et al., 2016). It is made fully available with a unique digital object identifier, in a CSV format that can be processed by most widely used tools, and is released under an Apache 2.0 license. Since the textual content of news articles may be copyrighted, we do not include them in the dataset, but instead provide a software (https://doi.org/10.5281/zenodo.10492913) to allow researchers to collect the content themselves.